OpenStudy (anonymous):

The polynomial x 3 + 5x 2 -57x -189 expresses the volume, in cubic inches, of a shipping box, and the width is (x+3) in. If the width of the box is 15 in., what are the other two dimensions? ( Hint: The height is greater than the depth.) A. height: 19 in. depth: 5 in. B.height: 21 in. depth: 5 in. c.height: 19 in depth: 7 in. D. height: 21 in. depth: 7 in. its not c.

OpenStudy (anonymous):

@DatChinookGuy do you think you can help with this one too

OpenStudy (wolf1728):

If width = x+3 = 15 Then x = 12 Putting 12 into the equation x 3 + 5x 2 -57x -189 then

OpenStudy (wolf1728):

box is 1,575 cubic inches

OpenStudy (wolf1728):

The answer is B because 21 * 5 * 15 = 1,575

OpenStudy (anonymous):

so you would plug in 12 for all the x's

OpenStudy (wolf1728):

Yes

OpenStudy (anonymous):

okay after you find 1575 then what would you do

OpenStudy (wolf1728):

we know width = 15 So going by the multiple choice answers we have to see what the 2 numbers times each other times 15 equal 1,575

OpenStudy (anonymous):

If width = x+3 = 15 Then x = 12 Putting 12 into the equation x 3 + 5x 2 -57x -189 then conduct the box which is 1,575 cubic inches. Then your answer is B because 21 * 5 * 15 = 1,575 However when doing this you MUST plug in 12 for all of the x's to signal the equation correctly and easier. Hope this helps, have a nice day :) In conclusion as said, B is your answer: \[height: 21 inches\] ===================== After you find 1575 you will take it down/divide it. \[Have a nice day \]

OpenStudy (anonymous):

oh okay i get it now thanks

OpenStudy (wolf1728):

u r welcome

OpenStudy (anonymous):

okay i have one more question

OpenStudy (anonymous):

No problem have a nice day, fire away with that last question! :)

OpenStudy (wolf1728):

Okay go ahead

OpenStudy (anonymous):

What are the zeros of the function? Graph the function. y=(x+3)(x-3)(x-4)

OpenStudy (wolf1728):

-3 +3 and +4

OpenStudy (wolf1728):

And you need a graph right?

OpenStudy (wolf1728):

The equation is X^3 -4x^2 -9x + 36

OpenStudy (anonymous):

i figured out the graph on my own